# Disentangling Representations through Multi-task Learning

### v2

@@ A.1 Definitions
In the main text we used the terms "abstractness" and "disentanglement". Since there exists some ambivalence about their meaning in the literature, we would like to strictly define them here. We will be using definitions adapted from Ostojic and Fusi [59]:
  • Abstract representation of $x_{1},\dots,x_{n}$ represents each $x_{i}$ linearly and approximately mutually orthogonally. Thus, abstractness ensures a simple linear map can decode each $x_{i}$ regardless of variation in $x_{j\neq i}$.
@@
Note that under this definition, axis-alignment is not a requirement for disentanglement (also see Higgins et al. [31]). Our work suggests that the computer science and neuroscience communities should adopt this broader definition of disentanglement, because otherwise we might be missing cases where the factors are not axis-aligned, but they are still orthogonal and can still be isolated by a linear decoder. Our argument is that there is nothing special about individual factors being encoded by individual neurons. Rather, we think that allowing for mixed representations within the definition of disentanglement leads to a more holistic view of disentanglement. A contribution of this work, along with others [37], is to bring this argument to the forefront.
@@ A.4 Derivation of theoretical $r^{2}$ for optimal multi-task classifiers
A.5 GPT-2 experiments
@@
A.6 Finding fixed points and linearization of dynamics
To find approximate fixed points of RNN dynamics after training, we follow a standard procedure outlined in Sussillo and Barak [70]. Specifically, we keep network weights fixed, provide no inputs to the network, and instead optimize over hidden activity. Specifically, we penalize any changes in the hidden activity, motivating the network to find stable states of the dynamics in the absence of input, i.e. attractors of the dynamics. This process finds all states of accumulated evidence that can be stored in this network as short-term memory. Network dynamics could then leverage these states to maintain and update the internal representation of the ground truth $\mathbf{x}^{*}$ on a single trial level, and drive downstream decisions.
Then for every approximate fixed point $\mathbf{z}^{f}$, we linearize RNN dynamics around it and estimate the eigenmodes which describe how the system behaves in a small region $\delta\mathbf{z}$ around $\mathbf{z}^{f}$. Specifically, following Sussillo and Barak [70], Mante et al. [49] we take the difference system $\delta\mathbf{z}(t+t_{0})=\mathbf{z}(t+t_{0})-\mathbf{z}(t_{0})$ and linearize it, i.e.
@@ A.6 Finding fixed points and linearization of dynamics
where $\dot{\mathbf{z}}=\mathbf{F}(\mathbf{z})$ is the function describing the RNN dynamics and $\mathbf{M}\equiv\mathbf{F}^{\prime}(\mathbf{z}^{f})$ is its Jacobian at $\mathbf{z}^{f}$. To estimate $\mathbf{F}^{\prime}(\mathbf{z}^{f})$, we let network dynamics run in the absence of inputs for one time step $\Delta t$ starting from $\mathbf{z}^{f}$, i.e. $\delta\mathbf{z}(\Delta t)=\mathbf{z}(\Delta t)-\mathbf{z}^{f}$, and autodifferentiate $\delta\mathbf{z}(\Delta t)$. We then perform eigendecomposition of $\mathbf{M}$ and report the eigenvalues around each approximate fixed point. Eigenvalues near 00 indicate that the difference system $\delta\mathbf{z}(t)=\mathbf{z}(t)-\mathbf{z}^{f}$ changes slowly over time, i.e. they correspond to "slow" dimensions in network dynamics which can integrate inputs and maintain them over time (continuous attractors) [3, 49].
A.7 State-space efficiency of context-dependent computation
The workhorse model for computational neuroscience has been context-dependent computation, where tasks are carried out one at a time and task identity is cued to the RNN by a one-hot vector [49]. However, this approach can be algorithmically inefficient, because as we show here it scales linearly with the number of tasks $N_{\textit{task}}$, and exponentially with input dimensionality $D$. This is because context-dependent computation utilizes different parts of the state space for different tasks, and the resulting representations collapse to what is minimally required for each task (also see [49, 77]). This can be detrimental for brains, which need to pack a lot of computation within a large yet limited neural substrate. In contrast, abstract representations are general, compact [48], can be used for any downstream task involving the same variables, scale linearly with $D$, and as we show readily emerge from relatively simple tasks.
To compare context-dependent decision making, where one task is performed at a time [49], to multitasking, in terms of state-space usage efficiency, we train RNNs to perform context-dependent decisions on the same tasks encountered in the main text. Compared to the network in the main text (Figure 2b), the RNN now also receives a one-hot task rule vector indicating the current task, and it outputs the decision for that task only (Figure S1b). We have also omitted the non-linear encoder, making the tasks easier. We train the RNN for two tasks, one task at a time, in interleaved batches (Figure S1a). In one task, the RNN is required to decide which stream has more evidence, and at the other whether the sum of evidence across streams exceeds a certain decision threshold (here 0).
@@ A.7 State-space efficiency of context-dependent computation
We find that in this setting the network is not incentivized to learn abstract representations. Instead a separate line attractor is present in the dynamics for each task (red and orange x’s in Figure S1c); one of them is presumably tracking the difference of evidence (similar to Yang and Wang [77] but for independent evidence streams) and the other the sum of evidence. That is to say, the task rule biases the network to learn different computations in separate regions of state space, as in Mante et al. [49]. As a result, the 2D latent space has collapsed and cannot be decoded from network activity; therefore generalization to any task that involves these two variables is not possible.
@@
Overall, we believe that multitasking may present a paradigm swift for generalizable representation learning in biological and artificial systems alike. That is not to say that context-dependent representations are not useful; they are great at leveraging the compositional structure of tasks [78, 20], but tend to overfit to the specifics of the task, while multitask representations serve as world models applicable to various scenarios. Both types of representations are likely to be found in the brain.
A.8 Robustness to other noise distributions and correlated inputs
@@
Figure S2: Disentanglement and factor correlations. (a) We introduce strong correlations in the latent factors, rendering parts of the state space virtually invisible to the network during pre-training (trained for a total of 24 classification tasks). (b) Despite that, generalization performance is excellent for correlations very close to 1. Once the factors are perfectly, performance drops significantly. This implies that the network can learn an abstract representation from correlated inputs, as long as there is some signal about the factors independently. This finding goes beyond [37] to show that the multi-task learning setting allows OOD generalization when the distribution during training the RNN itself is vastly different that the one during testing.
Finally, our theory pointed out sigmoid functions as a choice for activation function because of their close resemblance to the Gaussian CDF, resulting in the best OOD $r^{2}$. Still we find that for an asymmetric noise distribution (Gumbel) whose CDF does not match sigmoid functions well, there is only a slight drop in performance (median $r^{2}=0.95$ from $0.96$). Therefore the conditions for the activation function/CDF should be quite lax; any monotonic bijective function should work with small performance drop. This drop in performance is because the representation would be “stretched out” and “compressed” in a non-linear manner in regions where there is discrepancy between the noise CDF and the activation function. But this nonlinear squishing (determined by the term $\Phi^{-1}(g(\mathbf{Z}(t)))$) would be geometrically inoffensive — no cutting or gluing together would be required to map from $\mathbf{Z}(t)$ to a linear representation of $\mu(t)$. As a result, the representations would remain approximately linearly decodable.
A.9 Nonlinear classification boundaries and interleaved learning
@@
For that we perform interleaved training of both tasks (i.e. train in batches sampled from one of the tasks at a time), a setting where neural networks excel at, compared to humans who excel at blocked training, where tasks are learned sequentially (but see Flesch et al. [25]). Figure S3b shows that the network still learns a disentanlged, two-dimensional continuous attractor. OOD generalization for this network is excellent, and almost identical to ID performance (median $r^{2}=0.94,0.97$ respectively). Overall, we conclude that our framework extends to interleaved learning of a mixture of linear and nonlinear boundaries, which better reflects the challenges encountered by agents in the real world. Note that during interleaved training, linear and non-linear tasks are not performed simultaneously; yet they are in immediate succession which can also place pressure to the network to gradually learn representations that satisfy all tasks. The relation between multi-task and interleaved learning is a promising topic for future research.
A.10 Abstract representations are learned for a free reaction time, integrate to bound task
In the main text we trained networks to produce a response at the end of the trial. However, in many situations agents are free to make a decision whenever they are certain enough. Therefore, we here seek to extend our framework to free reaction time (RT) decisions. A canonical model accounting for choices and reaction times in humans and animals is the drift-diffusion model [40, 11]. It is composed of an accumulator that integrates noisy evidence over time, until a certain amount of certainty, represented by a bound, is reached, triggering a decision. In the linear classification task setting, the accumulated amount of evidence at time $t$ for a line with slope $\alpha$, $A_{\alpha}(t)$ is given by:
@@ A.10 Abstract representations are learned for a free reaction time, integrate to bound task
Intuitively, $A_{\alpha}(t)$ reflects the amount of confidence at time $t$ that the ground truth $\mathbf{x}^{*}$ lies above or below the classification line with slope $\alpha$. Essentially, the network has to explicitly report distance from the classification lines, not just in which side of the line $\mathbf{x}^{*}$ lies for that trial. We set the decision bound to ±$\pm 5$, and plot the accumulators $A_{\alpha}$ for all lines in Figure S4a. Note that once the bound is reached a decision is effectively made and $A_{\alpha}$ is kept constant. Also, instead of using continuous values, we quantize $A_{\alpha}$, because it is going to be used as target signal to train the network, and we do not want to introduce a strong inductive bias towards integrating the evidence streams.
@@
Figure S4b shows that in this setting the network still learns a two-dimensional continuous attractor of the latent space. Furthermore, the free RT outperforms the fixed RT network from the main text (Figure S4c) for a small number of tasks, since it is explicitly required to report distance from the classification lines. However, as our theory shows (Lemma B.3)) the fixed RT network is also implicitly reporting distance from the boundaries, when behaving like an optimal multi-task classifier, which explains the similar performance for a larger number of tasks. Overall, we showed that our setting accounts for naturalistic free RT decisions, and provides theoretical justification for the importance of confidence signals in the brain [66, 51].
A.11 Relation to neuroscience literature
@@
An ongoing debate in the brain sciences is whether to solve tasks the brain learns abstracts representations, or simple input-output mappings. Here we show that training RNNs to multitask results in shared, disentangled representations of the latent variables, in the form of continuous attractors. In this multitask setting, one task acts as a regularizer for the others, by not letting the representation collapse, or overfit, to specific tasks [79].
@@
Our findings directly link to two important neuroscientific findings: spatial cognition and value-based decision-making. First, the tasks here bear close resemblance to path-integration, i.e. the ability of animals to navigate space only relying on their proprioceptive sense of linear and angular velocity [54, 12, 73]. In path-integration animals integrate velocity signals to get location, while here we integrate noisy evidence to get rid of the noise. In path-integration, networks have to explicitly report distances, while in our setting distances are estimated implicitly (Lemma B.3)). We learn abstract representations in the form of a 2D "sheet" continuous attractor, while the computational substrate for path integration is a 2D toroidal attractor [26, 69] – not an abstract representation. The conditions under which a 2D sheet vs. toroidal continuous attractor is learned is a potential area of future research. Second, decision making experiments in monkeys result in a 2D abstract representation in the medial frontal cortex, which supports novel inferential decisions [6]. Likewise, context-dependent decision-making experiments in humans also resulted in orthogonal, abstract representations [25].

### v3

@@ A.1 Definitions
In the main text we used the terms ”abstractness” and ”disentanglement”. Since there exists some ambivalence about their meaning in the literature, we would like to strictly define them here. We will be using definitions adapted from Ostojic & Fusi (2024):
  • An abstract representation of latent factors $x_{1},\dots,x_{n}$ represents each $x_{i}$ linearly and approximately mutually orthogonally. Thus, abstractness ensures a simple linear map can decode each $x_{i}$ regardless of variation in $x_{j\neq i}$.
@@
Note that under this definition, axis-alignment is not a requirement for disentanglement (also see Higgins et al. (2018)). Our work suggests that the computer science and neuroscience communities should adopt this broader definition of disentanglement, because otherwise we might be missing cases where the factors are not axis-aligned, but they are still orthogonal and can still be isolated by a linear decoder. Our argument is that there is nothing special about individual factors being encoded by individual neurons. Rather, we think that allowing for mixed representations within the definition of disentanglement leads to a more holistic view of disentanglement. A contribution of this work, along with others (Johnston & Fusi, 2023), is to bring this argument to the forefront.
@@ A.4 Derivation of theoretical $r^{2}$ for optimal multi-task classifiers
A.5 More nonlinear encoding
Figure S1: OOD generalization is robust to choice of encoder nonlinearity.
In the main text we used ReLU nonlinearities for the encoder $f$. Here we extend our findings to more nonlinear observation maps which are likely to be encountered in the real world, e.g. ones with power-law nonlinearities. Recent work showed that the choice of activation function can influence the geometry of representations (Alleman et al., 2024). Interestingly, we find that replacing ReLUs in the encoder with a quadratic nonlinearity results in virtually identical OOD generalization performance compared to Figure 3a (Figure S1). Therefore we conclude that our setting is robust to the choice of encoder nonlinearity, even when the nonlinearity is not injective, going beyond our theoretical proofs (Section B.3).
A.6 GPT-2 experiments
@@
A.7 Finding fixed points and linearization of dynamics
To find approximate fixed points of RNN dynamics after training, we follow a standard procedure outlined in Sussillo & Barak (2013). Specifically, we keep network weights fixed, provide no inputs to the network, and instead optimize over hidden activity. Specifically, we penalize any changes in the hidden activity, motivating the network to find stable states of the dynamics in the absence of input, i.e. attractors of the dynamics. This process finds all states of accumulated evidence that can be stored in this network as short-term memory. Network dynamics could then leverage these states to maintain and update the internal representation of the ground truth $\mathbf{x}^{*}$ on a single trial level, and drive downstream decisions.
Then for every approximate fixed point $\mathbf{z}^{f}$, we linearize RNN dynamics around it and estimate the eigenmodes which describe how the system behaves in a small region $\delta\mathbf{z}$ around $\mathbf{z}^{f}$. Specifically, following Sussillo & Barak (2013); Mante et al. (2013) we take the difference system $\delta\mathbf{z}(t+t_{0})=\mathbf{z}(t+t_{0})-\mathbf{z}(t_{0})$ and linearize it, i.e.
@@ A.7 Finding fixed points and linearization of dynamics
where $\dot{\mathbf{z}}=\mathbf{F}(\mathbf{z})$ is the function describing the RNN dynamics and $\mathbf{M}\equiv\mathbf{F}^{\prime}(\mathbf{z}^{f})$ is its Jacobian at $\mathbf{z}^{f}$. To estimate $\mathbf{F}^{\prime}(\mathbf{z}^{f})$, we let network dynamics run in the absence of inputs for one time step $\Delta t$ starting from $\mathbf{z}^{f}$, i.e. $\delta\mathbf{z}(\Delta t)=\mathbf{z}(\Delta t)-\mathbf{z}^{f}$, and autodifferentiate $\delta\mathbf{z}(\Delta t)$. We then perform eigendecomposition of $\mathbf{M}$ and report the eigenvalues around each approximate fixed point. Eigenvalues near 00 indicate that the difference system $\delta\mathbf{z}(t)=\mathbf{z}(t)-\mathbf{z}^{f}$ changes slowly over time, i.e. they correspond to ”slow” dimensions in network dynamics which can integrate inputs and maintain them over time (continuous attractors) (Amari, 1977; Mante et al., 2013).
A.8 State-space efficiency of context-dependent computation
The workhorse model for computational neuroscience has traditionally been context-dependent computation, where tasks are carried out one at a time and task identity is cued to the RNN by a one-hot vector Mante et al. (2013). However, this approach can be algorithmically inefficient, because as we show here it scales linearly with the number of tasks $N_{\textit{task}}$, and exponentially with input dimensionality $D$. This is because context-dependent computation utilizes different parts of the state space for different tasks, and the resulting representations collapse to what is minimally required for each task (also see (Mante et al., 2013; Yang & Wang, 2020)). This can be detrimental for brains, which need to pack a lot of computation within a large yet limited neural substrate. In contrast, abstract representations are general, compact (Ma et al., 2022), can be used for any downstream task involving the same variables, scale linearly with $D$, and as we show readily emerge from relatively simple tasks.
To compare context-dependent decision making, where one task is performed at a time (Mante et al., 2013), to multitasking, in terms of state-space usage efficiency, we train RNNs to perform context-dependent decisions on the same tasks encountered in the main text. Compared to the network in the main text (Figure 2b), the RNN now also receives a one-hot task rule vector indicating the current task, and it outputs the decision for that task only (Figure S2b). We have also omitted the non-linear encoder, making the tasks easier. We train the RNN for two tasks, one task at a time, in interleaved batches (Figure S2a). In one task, the RNN is required to decide which stream has more evidence, and at the other whether the sum of evidence across streams exceeds a certain decision threshold (here 0).
@@ A.8 State-space efficiency of context-dependent computation
We find that in this setting the network is not incentivized to learn abstract representations. Instead a separate line attractor is present in the dynamics for each task (red and orange x’s in Figure S2c); one of them is presumably tracking the difference of evidence (similar to Yang & Wang (2020) but for independent evidence streams) and the other the sum of evidence. That is to say, the task rule biases the network to learn different computations in separate regions of state space, as in Mante et al. (2013). As a result, the 2D latent space has collapsed and cannot be decoded from network activity; therefore generalization to any task that involves these two variables is not possible.
@@
Overall, we believe that multitasking may present a paradigm swift for generalizable representation learning in biological and artificial systems alike. That is not to say that context-dependent representations are not useful; they are great at leveraging the compositional structure of tasks (Yang et al., 2019; Driscoll et al., 2022), but tend to overfit to the specifics of the task, while multitask representations serve as world models applicable to various scenarios. Both types of representations are likely to be found in the brain. One possibility is that context-dependent representations may emerge as a first quick solution to a task, while disentangled representations come about with more experience or when more tasks are performed over time to support better generalization.
A.9 Robustness to other noise distributions and correlated inputs
@@
Figure S3: Disentanglement and factor correlations. (a) We introduce strong correlations in the latent factors, rendering parts of the state space virtually invisible to the network during pre-training (trained for a total of 24 classification tasks). (b) Despite that, generalization performance is excellent for correlations very close to 1. Once the factors are perfectly correlated, performance drops significantly. This implies that the network can learn an abstract representation from correlated inputs, as long as there is some signal about the factors independently. This finding goes beyond (Johnston & Fusi, 2023) to show that the multi-task learning setting allows OOD generalization when the distribution during training the RNN itself is vastly different that the one during testing.
Finally, our theory pointed out sigmoid functions as a choice for activation function because of their close resemblance to the Gaussian CDF, resulting in the best OOD $r^{2}$. Still we find that for an asymmetric noise distribution (Gumbel) whose CDF does not match sigmoid functions well, there is only a slight drop in performance (median $r^{2}=0.95$ from $0.96$). Therefore the conditions for the activation function/CDF should be quite lax; any monotonic bijective function should work with small performance drop. This drop in performance is because the representation would be “stretched out” and “compressed” in a non-linear manner in regions where there is discrepancy between the noise CDF and the activation function. But this nonlinear squishing (determined by the term $\Phi^{-1}(g(\mathbf{Z}(t)))$) would be geometrically inoffensive — no cutting or gluing together would be required to map from $\mathbf{Z}(t)$ to a linear representation of $\mu(t)$. As a result, the representations would remain approximately linearly decodable. Monotonicity and bijectivity are quite mild assumptions for the activation function used by neurons in the brain.
A.10 Nonlinear classification boundaries and interleaved learning
@@
For that we perform interleaved training of both tasks (i.e. train in batches sampled from one of the tasks at a time), a setting where neural networks excel at, compared to humans who excel at blocked training, where tasks are learned sequentially (but see Flesch et al. (2022)). Figure S4b shows that the network still learns an abstract, two-dimensional continuous attractor. OOD generalization for this network is excellent, and almost identical to ID performance (median $r^{2}=0.94,0.97$ respectively). Overall, we conclude that our framework extends to interleaved learning of a mixture of linear and nonlinear boundaries, which better reflects the challenges encountered by agents in the real world. Note that during interleaved training, linear and non-linear tasks are not performed simultaneously; yet they are in immediate succession which can also place pressure to the network to gradually learn representations that satisfy all tasks. The relation between multi-task and interleaved learning is a promising topic for future research.
A.11 Abstract representations are learned for a free reaction time, integrate to bound task
In the main text we trained networks to produce a response at the end of the trial. However, in many situations agents are free to make a decision whenever they are certain enough. Therefore, we here seek to extend our framework to free reaction time (RT) decisions. A canonical model accounting for choices and reaction times in humans and animals is the drift-diffusion model (Krajbich et al., 2010; Brunton et al., 2013). It is composed of an accumulator that integrates noisy evidence over time, until a certain amount of certainty, represented by a bound, is reached, triggering a decision. In the linear classification task setting, the accumulated amount of evidence at time $t$ for a line with slope $\alpha$, $A_{\alpha}(t)$ is given by:
@@ A.11 Abstract representations are learned for a free reaction time, integrate to bound task
Intuitively, $A_{\alpha}(t)$ reflects the amount of confidence at time $t$ that the ground truth $\mathbf{x}^{*}$ lies above or below the classification line with slope $\alpha$. Essentially, the network has to explicitly report distance from the classification lines, not just in which side of the line $\mathbf{x}^{*}$ lies for that trial. We set the decision bound to ±$\pm 5$, and plot the accumulators $A_{\alpha}$ for all lines in Figure 2a. Note that once the bound is reached a decision is effectively made and $A_{\alpha}$ is kept constant. Also, instead of using continuous values, we quantize $A_{\alpha}$, because it is going to be used as target signal to train the network, and we do not want to introduce a strong inductive bias towards integrating the evidence streams.
@@
Figure S5b shows that in this setting the network still learns a two-dimensional continuous attractor of the latent space. Furthermore, the free RT outperforms the fixed RT network from the main text (Figure S5c) for a small number of tasks, since it is explicitly required to report distance from the classification lines. However, as our theory shows (Lemma B.3)) the fixed RT network is also implicitly reporting distance from the boundaries, when behaving like an optimal multi-task classifier, which explains the similar performance for a larger number of tasks. Overall, we showed that our setting accounts for naturalistic free RT decisions, and provides theoretical justification for the importance of confidence signals in the brain (Rutishauser et al., 2018; Masset et al., 2020).
The importance of the confidence (i.e., calibrated likelihoods) of a network’s output, is a recurring theme in machine learning too (e.g. knowledge distillation (Bhargava et al., 2024)). We here show that confidence fundamentally connects to how neural networks construct world models, either directly (integrate-to-bound task here) or indirectly (classification tasks in main text). Under this framework, knowledge distillation can be cast as smaller models directly copying the world models (logits) of larger ones.
A.12 Quantification of sparsity
In the main text, we observed that RNN representations are sparse. We here seek to more precisely quantify the sparsity in these networks, and investigate how it is affected by the number of tasks $N_{\textit{task}}$, latent dimensionality $D$, and specific recurrent architecture. To do so, we sample $n=1000$ ground truth vectors $\mathbf{x}^{*}$ randomly for every network, and compute the sparseness (Vinje & Gallant, 2000) of a neuron in the hidden layer as:
$S=\frac{1-\left(\frac{\sum\left(z_{i}/n\right)^{2}}{\sum\left(z_{i}^{2}/n\right)}\right)}{1-\frac{1}{n}}*100\,\%$ (10)
where $z_{i}$ is the steady-state response of the neuron to ground-truth stimulus $i$. Sparseness ranges from 00 to $100\,\%$, with greater sparseness indicating greater selectivity of the neuron to stimuli. Then, the sparsity of a network is given as the average of the sparseness of all its neurons.
Figure S6: Quantification of sparsity as a function of $N_{\textit{task}}$, $D$ and recurrent architecture choice. (a) Sparsity of a recurrent network as a function of number of tasks and network architecture. Five networks trained for each network configuration. Greater levels of sparsity indicate that the network activations are more sparse. (b) Sparsity of RNNs as a function of number of tasks and latent dimensionality $D$. Five network are trained for each combination of $(N_{\text{task}},D)$.
Figure S6a shows that RNNs and non-leaky RNNs are very sparse, with sparsity values around $90\,\%$ for different values of $N_{\textit{task}}$, supporting the claim in the main text. LSTMs on the other hand, which are less brain-like [^4] LSTMs architecturally enforce intricate, high-capacity multiplicative gating mechanisms, while in biological neural networks gating has to be learned. For other aspects of biological implausibility of LSTMs compared to RNNs, see Appendix B in (Soo et al., 2023)., have lower sparsity values, although interestingly sparsity increases with $N_{\textit{task}}$. Notably, we did not do anything to promote sparsity (e.g. regularization) in these networks. Therefore we conclude that sparsity naturally emerges from the optimization objective of multitask learning, particularly in architectures that are more brain-like.
Next we wondered how latent space dimensionality $D$ would affect sparsity in our trained RNNs. Figure S6b shows that networks remain very sparse for the whole range of dimensionality $D$ tested in the main text, with sparsity values above $75\,\%$. Greater dimensionality results in less sparsity on average, which is expected since $N_{\textit{neu}}=64$ in our networks, therefore a significant amount of their capacity must be used as $D$ increases. This effect plays in only as $N_{\textit{task}}$ increases, as networks will only learn to disentangle the input dimensions that are spanned by the tasks, as our theory predicts. Overall, there seems to be a proportional relationship between the number of active neurons and dimensionality $D$, as long as there are enough tasks to uncover the $D$ latents.
A.13 Relation to neuroscience literature
@@
An ongoing debate in the brain sciences is whether to solve tasks the brain learns abstracts representations, or simple input-output mappings. Here we show that training RNNs to multitask results in shared, disentangled representations of the latent variables, in the form of continuous attractors. In this multitask setting, one task acts as a regularizer for the others, by not letting the representation collapse, or overfit, to specific tasks (Zhang & Yang, 2017).
@@
Our findings directly link to two important neuroscientific findings: spatial cognition and value-based decision-making. First, the tasks here bear close resemblance to path-integration, i.e. the ability of animals to navigate space only relying on their proprioceptive sense of linear and angular velocity (Mittelstaedt & Mittelstaedt, 1980; Burak & Fiete, 2009; Vafidis et al., 2022; Sorscher et al., 2023). In path-integration animals integrate velocity signals to get location, while here we integrate noisy evidence to get rid of the noise. In path-integration, networks have to explicitly report distances, while in our setting distances are estimated implicitly (Lemma B.3)). We learn abstract representations in the form of a 2D ”sheet” continuous attractor, while the computational substrate for path integration is a 2D toroidal attractor (Gardner et al., 2022; Sorscher et al., 2023) – not an abstract representation. The conditions under which a 2D sheet vs. toroidal continuous attractor is learned is a potential area of future research. Second, decision making experiments in monkeys result in a 2D abstract representation in the medial frontal cortex, which supports novel inferential decisions (Bongioanni et al., 2021). Likewise, context-dependent decision-making experiments in humans also resulted in orthogonal, abstract representations (Flesch et al., 2022).
A.14 Biological plausibility of multi-task learning
While our theory stems from parallel processing, i.e. multi-task learning, it is not contingent upon the parallel execution of multiple tasks, i.e. multitasking, or the receipt of rich supervisory feedback from the environment in parallel. Behaviorally, the agent need only perform one action, the one most appropriate to its current internal state (e.g. its level of thirst vs. hunger might control the slope of the decision boundary in the 2D latent space of water & food). What we posit is that tasks that have been performed by the agent before and rely on the same input are still resolved somewhere in the brain, by the brain circuits (e.g. cortical columns Hawkins et al. (2019)) previously responsible for them, instead of the entire decision-making brain area focusing only on the current task (Mante et al., 2013). Therefore, the output of these tasks is still placing pressure on the representation, even though they are not actively driving behavior. In other words, our theory assumes competence at $N_{\text{task}}$ tasks, independently of when and how that competence was achieved. We feel that this is a more natural way of thinking about how the brain manages different tasks, with older tasks still leaving traces somewhere in the brain (Losey et al., 2024); after all, biological agents are remarkable because they achieve high performance on many tasks. This theory is also closely related to the widely observed phenomenon of memory replay (Foster & Wilson, 2006), or mental simulation of counterfactuals (Jensen et al., 2024). A future direction to further enhance the biological relevance of our work would be to investigate the relation between multi-task learning and slow, interleaved learning (see Section A.10), in a continual learning setting.
A.15 Limitations and future directions
A limitation of the present work is that factorization is assumed. Yet not all problems are factorizable, or should be factorized. For instance, a more coarse-grained understanding of the world, that doesn’t disentangle all factors, might be more suitable in many cases, and that might be reflected in the nature of the tasks. Furthermore, we focus on canonical cognitive neuroscience tasks which are somewhat removed from standard ML benchmarks. Normally, disentanglement methods would be tested against a benchmark such as dSprites (Matthey et al., 2017); however to the best of our knowledge no such benchmark exists for sequential tasks where evidence has to be aggregated over time. Future work could endeavor to apply our setting to richer tasks, like extracting latent item attributes from item embeddings when sequential decisions are made in online retailer settings.
Our theory is agnostic to the way by which competence at multiple tasks is achieved. Thus, a natural next step is to investigate whether disentangled representations exist in a wider range of models capable of solving multiple tasks. A prime example is large language models that display excellent zero- and few-shot generalization capabilities, with progress already made in that direction (Templeton et al., 2024). Moreover, the pre-training objective for LLMs (cross entropy loss/likelihood maximization) fits well within our theoretical framing on (approximately) optimal multi-classifiers. Another application area, as already mentioned, is neuroscience; animals are naturally competent at multiple tasks, thus our work provides theoretical justification for why disentangled representations have been found in many brain areas, and motivates looking for more.
Our experiments showed parsimony of our theoretical results under conditions not covered by our theory, including non-injective observation maps (Section A.10) and decision boundaries (Section A.11) which is encouraging for testing our findings on settings beyond what is strictly covered by the theory. It would be interesting to see how the theoretical insights generalize to different task geometries, for example those implied by self-supervised learning applications (e.g. image patch-filling, next-token prediction, iterative de-noising). The connection between our framework and self-supervised learning is deep and promising. Both frameworks share a common structure, where an underlying latent truth (e.g., objects in an image and their relationships) is inferred. Each objective (e.g., filling in missing image patches) contributes synergistically to understanding the latent space as a whole. A similar logic applies to predicting words, where the latent “meaning” of a sentence is shared, whether in a causal (e.g., LLMs) or masked setting (e.g., BERT). Our study is a first effort towards understanding such parallel learning, and providing guarantees for its performance.